OpenStudy (anonymous):

MEDAL FOR THE BEST ANSWER -5 1/2 + 6 3/4 + (-4 1/4 ) = _______ A. 16 1/2 B. -3 C. 8 D. 5 1/2

OpenStudy (anonymous):

First, Make the mixed numbers into improper fractions. Do you know how to do that?

OpenStudy (anonymous):

To turn a mixed number into an improper fraction, you use this method: \[6\frac{ 3 }{ 4 } = 6 \times 4 = 24 + 3 = 27\] Then, you put that number over the same denominator (bottom number of the fraction): \[\frac{ 27 }{ 4 }\]

OpenStudy (anonymous):

I don't get sorry I'm confused what to do with 27/4

OpenStudy (anonymous):

I made the mixed number 6 3/4 into 27/4....now, you need to turn the other mixed number in the equation into improper fractions using the same method. The method is: 1. Multiply the whole number by the denominator (bottom number of fraction). 2. Add the numerator (top number of fraction) to that number and put it over the same denominator. FOR EXAMPLE - this is how you would change a mixed number into an improper fraction: For example, 1 2/3: 1*3 = 3 3 + 2 = 5 Final answer:5/3 ^ That was just an example on how to do it.

OpenStudy (anonymous):

Ok I understand that, but what about negative number like -5 1/2

OpenStudy (anonymous):

It would be: \[-5 \times 2 + 1 = ?\] Then put that number over the original denominator of the fraction (2)

OpenStudy (anonymous):

Tell me what you get.